### v2

# MoTCoder: Elevating Large Language Models with Modular of Thought for Challenging Programming Tasks

Abstract
Large Language Models (LLMs) have showcased impressive capabilities in handling straightforward programming tasks. However, their performance tends to falter when confronted with more challenging programming problems. We observe that conventional models often generate solutions as monolithic code blocks, restricting their effectiveness in tackling intricate questions. To overcome this limitation, we present Modular-of-Thought Coder (MoTCoder). We introduce a pioneering framework for MoT instruction tuning, designed to promote the decomposition of tasks into logical sub-tasks and sub-modules. Our investigations reveal that, through the cultivation and utilization of sub-modules, MoTCoder significantly improves both the modularity and correctness of the generated solutions, leading to substantial relative pass@1 improvements of 12.9% on APPS and 9.43% on CodeContests. Our codes are available at https://github.com/dvlab-research/MoTCoder.
[…]
itemsep=0pt,partopsep=0pt,parsep=0pt,topsep=0pt
[…]
1 Introduction
It has been a longstanding objective in the field of artificial intelligence to develop systems capable of generating executable and functionally correct computer programs to address intricate problems (Manna & Waldinger, 1971). Recently, there has been considerable attention on Large Language Models (LLMs) (Brown et al., 2020; OpenAI, 2023; Chowdhery et al., 2022; Anil et al., 2023; Hoffmann et al., 2022; Rae et al., 2021; Zeng et al., 2022; Touvron et al., 2023; Zhang et al., 2022), which showcases remarkable success. These models leverage extensive pre-training and undergo further fine-tuning with detailed instruction data (Ouyang et al., 2022), leading to SOTA performance across diverse tasks. Initially designed for natural languages, these models have been expanded through the instruction fine-tuning for code modeling capabilities (Rozière et al., 2023; Black et al., 2021; Chen et al., 2021c). This extension has led to impressive performance in generating code from natural language problem descriptions (Li et al., 2023; Luo et al., 2023a; Wang et al., 2023a; Li et al., 2022b; Nijkamp et al., 2023; Zheng et al., 2023; Fried et al., 2022; Chen et al., 2021a; Wang et al., 2021).
Figure 1: APPS test results by pass@k (%). Comparison of our MoTCoder with previous SOTAs.
Figure 2: The illustration of our MoTCoder training framework. We first perform the modular-of-thought instructions evolution, including the filtration of instructions that do not undergo successful evolution. Subsequently, we delve into the MoT instruction tuning and testing of the model.
Figure 3: To evolve conventional instructions into modular-of-thought instructions, our MoT instruction evolution approach guides the models through a two-step process. Initially, the models are instructed to outline the necessary sub-modules, generating only their function headers and docstrings that describe their intended usage. The subsequent instruction then guides the model to implement these modules and integrate them into a cohesive final solution. This instruction is complemented by a one-shot example to further enhance understanding.
However, when confronted with highly intricate coding challenges, the current SOTA models still struggle to match the seasoned developers (Hendrycks et al., 2021b; Li et al., 2022a; Shinn et al., 2023), mainly due to their overly simplistic generation approach. Many previous LLMs produce the code solution as a single monolithic block, instead of breaking down of the task into logical sub-tasks. However, tackling highly complex problems with a single code module is practically unattainable. In contrast, adept developers devise modularized solutions, organizing them into high-level logical sub-modules. Subsequently, they complete and combine these components to efficiently enhance their final solutions.
In our Modular-of-Thought Coder (MoTCoder), our objective is to enhance the modularization capabilities of coding Large Language Models (LLMs) by introducing modular-of-thought prompting. This approach guides LLMs to break down their solutions into modular segments, where each segment represents an abstract function dedicated to a high-level logical sub-task. To train the model to adhere to the MoT prompt, we generate instructional data using a process termed MoT Code Instruction Evolution.
As depicted in Figure 2, our MoT code instruction evolution framework comprises two key stages: In the initial stage, we tailor the evolutionary prompt process. LLMs are instructed to outline necessary sub-modules, generating only their function headers and docstrings that describe their intended usage. Subsequently, the instruction guides the model in implementing these modules and eventually combining them into a comprehensive final solution. In the second stage, we fine-tune the LLM using our created code instruction-following training set, resulting in our MoTCoder model.
Figure 4: Illustration of modular code generation involves the model initially delineating the necessary sub-modules, each comprising a function header and a docstring detailing its intended purpose. Following this, the model proceeds to comprehensively implement each module in code and incorporates them as integral components of the finalized solution..
Our experiments reveal that MoTCoder significantly enhances LLM performance, establishing new SOTA results on challenging code tasks in APPS (Hendrycks et al., 2021b) and CodeContests  (Li et al., 2022a). Specifically, MoTCoder elevates the pass@1 performance by over 12.9% on APPS and 9.43% on CodeContests.
2 Related Works
2.1 Large Language Models
General LLMs. In recent times, Large Language Models (LLMs) have exhibited remarkable prowess across a wide array of tasks. Leading technology companies have made significant advancements in developing highly proficient closed-source LLMs, including OpenAI’s GPT3 and GPT4 (Brown et al., 2020; OpenAI, 2023), Google’s PaLM (Chowdhery et al., 2022; Anil et al., 2023), Bard [^1] https://bard.google.com/, DeepMind’s Chinchilla (Hoffmann et al., 2022), and Gopher (Rae et al., 2021), as well as Anthropic’s Claude [^2] https://www.anthropic.com/index/introducing-claude. The AI community has also observed the release of several open-source LLMs, where model weights are made publicly available. EleutherAI has contributed GPT-NeoX-20B (Black et al., 2022) and GPT-J-6B (Wang & Komatsuzaki, 2021). Google has released UL2-20B (Tay et al., 2022). Tsinghua University has introduced GLM-130B (Zeng et al., 2022). Meta has released OPT (Zhang et al., 2022) and LLaMA (Touvron et al., 2023).
Coding LLMs. Recent research has introduced a significant number of LLMs tailored for code-related tasks to address the challenges of code understanding and generation. Closed-source models include OpenAI’s Codex (Chen et al., 2021a) and Code-Davinci (Microsoft, 2023). Google has proposed PaLM-Coder (Chowdhery et al., 2022). These models excel on popular code completion benchmarks such as HumanEval (Chen et al., 2021b) and MBPP (Austin et al., 2021). On the open-source front, Salesforce has introduced CodeGen (Nijkamp et al., 2023), CodeT5 (Wang et al., 2021), and CodeT5+ (Wang et al., 2023b). Tsinghua University has contributed CodeGeeX (Zheng et al., 2023), and the BigCode Project has developed StarCoder (Li et al., 2023). These models have demonstrated significant advancements in code-related tasks.
2.2 Instruction Fine-Tuning
General Instruction Tuning. In its early stages, the core aim of instruction fine-tuning was to amplify the cross-task generalization capabilities of Language Models (LMs). This was accomplished by subjecting LMs to fine-tuning using an extensive corpus of public Natural Language Processing (NLP) tasks. Pioneering this approach, T5 (Raffel et al., 2020) underwent training on a diverse set of supervised text-to-text tasks. Subsequent endeavors like FLAN (Wei et al., 2022a), ExT5 (Aribandi et al., 2022), T0 (Sanh et al., 2022), and UnifiedQA (Khashabi et al., 2020) broadened the spectrum of tasks, fortifying the overall generalization capability of LMs. Noteworthy contributions from ZeroPrompt (Xu et al., 2022) and FLAN-T5 (Chung et al., 2022) pushed boundaries by incorporating thousands of tasks into their training pipelines. OpenAI has taken an alternative route by enlisting human annotators to contribute an extensive corpus of human instructions, encompassing diverse formats and a broad spectrum of task types. Building upon this dataset, OpenAI trained its GPT-3 (Brown et al., 2020) model to create InstructGPT (Ouyang et al., 2022), which better aligns with users’ inputs. This developmental trajectory has given rise to notable works such as ChatGPT. In the open-source realm, Alpaca (Taori et al., 2023) adopts the self-instruct method (Wang et al., 2022), leveraging ChatGPT to generate data for training. Vicuna (Chiang et al., 2023) utilizes user-shared conversations collected from ShareGPT.com to train its models. Introducing the Evol-Instruct method, WizardLM (Xu et al., 2023) involves evolving existing instruction data to generate more intricate and diverse datasets. In contrast to these generalized instruction fine-tuning approaches, WizardCoder (Luo et al., 2023b) applies the Evol-Instruct method specifically in the domain of Code LLMs.
Chain-of-Thought Instruction Tuning. While large language models have demonstrated impressive capabilities in handling straightforward programming tasks, their performance tends to diminish when faced with more challenging programming problems. A notable limitation is observed in conventional models, where solutions are often generated as monolithic code blocks, limiting their effectiveness in addressing intricate questions. Drawing inspiration from the success of chain-of-thought (CoT) prompting in enhancing performance in reasoning-based tasks (Zhou et al., 2023; Wei et al., 2022b; Kojima et al., 2022; Le et al., 2023), we introduce a pioneering framework for Modular-of-Thought (MoT) instruction tuning. MoT is designed to facilitate the decomposition of tasks into logical sub-tasks and sub-modules. Our investigations reveal that, through the cultivation and utilization of sub-modules, MoTCoder significantly enhances both the modularity and correctness of the generated solutions.
3 Methods
The depiction of our MoTCoder training framework is presented in  Figure 2. In the initial stage, we introduce the modular-of-thought instructions evolution and evolution examination in  Section 3.1. Following this, we elaborate on the MoT instruction tuning and testing of the model in  Section 3.2.
3.1 Modular-of-Thought Instruction Evolution
In this section, we propose the modular-of-thought instruct evolution approach. We initiate the evolution process with a given initial instruction dataset $D^{(0)}={(I_{n}^{(0)},\hat{O}_{n}^{(0)})}_{1\leq n\leq N}$, where $I_{n}^{(0)}$ represents the $n$-th instruction in $D^{(0)}$, $O_{n}^{(0)}$ is the corresponding output, and $N$ is the number of samples in $D^{(0)}$. In each evolution step, we upgrade all $I^{(k)}$ in $D^{(k)}$ to $I^{(k+1)}$ by employing an LLM instruction evolution prompt. Subsequently, we use the LLM to generate corresponding outputs $\hat{O}^{(k+1)}$ for the newly evolved $I^{(k+1)}$. Thus, we obtain an evolved instruction dataset $\{D^{(0)}\dots D^{(k+1)}\}$. By iteratively conducting $K$ evolutions, we sequentially acquire the final MoT instruction dataset $\{D^{(0)}\dots D^{(K)}\}$.
Our process for the evolution of modular-of-thought instructions comprises two key stages: MoT instructions evolution, and evolution examination, involving the filtration of instructions that do not successfully evolve. We commence by populating the instruction pool with the provided initial instruction dataset $D^{(0)}$. In each subsequent evolution epoch, normal instructions from the preceding epoch are retrieved from the pool. Subsequently, we employ the instruction evolver to advance each retrieved instruction and the instruction eliminator to evaluate whether the evolution is unsuccessful. Instructions that evolve successfully are integrated into the final pool, while unsuccessful ones are returned as they are, with the anticipation of achieving successful enhancement in the following evolution epoch.
Normal Instruction.
In general, a code sequence generated by a language model $\theta$ through the autoregressive sampling of tokens $\hat{o}_{t}$ is from the parameterized conditional distribution:
$\hat{o}_{t}\sim p_{\theta}(.|\hat{o}_{1:t-1},I),$ (1)
where $I$ represents the input instruction and $\hat{o}_{t}$ is the $t$-th of the flattened output sequence.
Modular-of-Thought Instruction.
Referring to the recent works in code revision (Le et al., 2023), our proposed methodology aims to evolute normal instructions into a sequential code generation process by leading the models through a two-step procedure.
  1. 1.
    Sub-modules. Initially, the models are instructed to outline the required sub-modules, generating only their function headers and docstrings describing their intended usage.
    $\hat{S_{i}}\sim p_{\theta}(.|\hat{S}_{1:i-1},I),$ (2)
    where $\hat{S}_{i}$ represents the $i$-th sub-module outlined by the model and $I$ represents the input instruction.
  2. 2.
    Final solution. The subsequent instruction guides the model to implement these modules and eventually combine them into a comprehensive final solution.
    $\hat{o}_{t}\sim p_{\theta}(.|\hat{o}_{1:t-1},\{\hat{S}_{i}\},I),$ (3)
    where $\hat{o}_{t}$ is the $t$-th of the flattened output sequence.
Modular-of-Thought Instruction Illustration.
The instruction is supplemented with a one-shot example, serving to prompt the model to adhere to the MoT instruction generation strategy. An illustration of the instruction prompt is presented in Figure 3 and further detailed in Figure 4. The instruction generated by MoTCoder encourages the model to decompose a program into sub-modules. This mirrors the methodology commonly employed by developers when addressing intricate coding tasks, where they systematically break down solutions into modular components.
(a) Modular Prompt
(b) Normal Prompt
Figure 5: Example of modular and normal code generation prompts. The modular instructions provide sub-modules including their function headers and docstrings outlining their intended usage. The final resolution executes these modules and then incorporates them into a comprehensive solution.
Evolution Examination.
We categorize the following scenarios as instances of instruction evolution failure:
  1. 1.
    The evolved instruction deviates from the modular-of-thought generation strategy, failing to follow the approach of initially generating sub-modules and then generating the main code.
  2. 2.
    During the first phase, namely the sub-module generation stage, either no sub-modules are generated, or global code is produced.
  3. 3.
    In the second phase, or the main code generation stage, there is either a lack of main code generation or the production of multiple main code segments.
Following the evolution process, we created a modular-of-thought instruction dataset comprising around 24k instructions. The dataset exhibits a distribution ratio of 45:55 between MoT and normal instruction data.
3.2 Modular-of-Thought Training and Inference
We present the MoT prompt training methodology through the introduction of a modular training prompt, as illustrated in Figure 5a. The instructions first specify sub-modules with their function headers and docstrings detailing their intended usage. The ultimate solution involves implementing these modules and subsequently integrating them into a final solution. The {input} in prompts is the problem description and the model is trained to complete the {response} in prompts. We conducted instruction tuning using the base model WizardCoder (Luo et al., 2023b) on our MoT instruction dataset for three epochs. This process took approximately 28 hours using 8 A800 GPUs.
Table 1: Experiments of modular ability. We conducted inference using both normal and modular prompts, as demonstrated in Figure 5 and show APPS test results by pass@1 (%). Comparable approaches include StarCoder (Li et al., 2023) and WizardCoder (Luo et al., 2023b) with the same size of 15B.
Model Size Prompt Introductory Interview Competition All
StarCoder 15B Normal 7.25 6.89 4.08 6.40
Modular 0.59 2.87 1.00 2.04
WizardCoder 15B Normal 26.04 4.21 0.81 7.90
Modular 0.39 0.46 0.66 0.81
Normal 33.80 19.70 11.09 20.80
MoTCoder 15B Modular 35.86 18.79 11.82 20.81
Table 2: Modular-of-thought instruction tuning ablation experiments. We compare the one-step and two-step modular-of-thought instruction tuning strategies. Models are inferred using both normal and modular prompts. APPs test results by pass@1 (%) are shown.
Model Size Prompt Introductory Interview Competition All
Normal 30.76 17.16 11.16 18.68
1-Step MoTCoder 15B Modular 31.69 17.37 11.47 19.05
Normal 33.80 19.70 11.09 20.80
2-Step MoTCoder 15B Modular 35.86 18.79 11.82 20.81
Table 3: Training Datasets. We compare MoTCoder training on only modular dataset and the mixture of modular and normal dataset. APPs test results by pass@1 (%) are shown.
Model Introductory Interview Competition All
Baseline 7.25 6.89 4.08 6.40
Modular Only 31.34 15.16 10.25 17.42
Mixture 33.80 19.70 11.09 20.80
Table 4: APPS test results by pass@k (%). Competitive open-source approaches include CodeT5 (Wang et al., 2021), fine-tuned GPT-Neo (Hendrycks et al., 2021a), GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), one-shot StarCoder (Li et al., 2023), and one-shot WizardCoder (Luo et al., 2023b). We also present results from closed-source models, namely text-davinci, code-davinci, and GPT3.5.
Model Size Pass@ Introductory Interview Competition All
CodeT5 770M 1 6.60 1.03 0.30 2.00
GPT-Neo 2.7B 1 14.68 9.85 6.54 10.15
5 19.89 13.19 9.90 13.87
GPT-2 0.1B 1 5.64 6.93 4.37 6.16
5 13.81 10.97 7.03 10.75
1.5B 1 7.40 9.11 5.05 7.96
5 16.86 13.84 9.01 13.48
GPT-3 175B 1 0.57 0.65 0.21 0.55
StarCoder 15B 1 7.25 6.89 4.08 6.40
WizardCoder 15B 1 26.04 4.21 0.81 7.90
MoTCoder 15B 1 33.80 19.70 11.09 20.80
text-davinci-002 - 1 - - - 7.48
code-davinci-002 - 1 29.30 6.40 2.50 10.20
GPT3.5 - 1 48.00 19.42 5.42 22.33
Table 5: APPS test results with code-revision by pass@1 (%). Revision indicates the rounds of revision after the first round of direct generation. Baselines include CodeT5 (Wang et al., 2021), text-davinci and WizardCoder (Luo et al., 2023b) and code-revision methods contain Self-edit (Zhang et al., 2023), CodeRL (Le et al., 2022), Self-repair (Olausson et al., 2023), and CodeChain (Le et al., 2023).
Model Size Revision Introductory Interview Competition All
CodeT5 770M - 6.60 1.03 0.30 2.00
CodeRL+CodeT5 770M 5 7.08 1.86 0.75 2.69
text-davinci-002 - - - - - 7.48
Self-edit+text-davinci-002 - 5 - - - 7.94
WizardCoder 15B - 26.04 4.21 0.81 7.90
CodeChain+WizardCoder 15B 5 26.29 7.49 3.75 10.50
MoTCoder 15B - 33.80 19.70 11.09 20.80
4 Experiments
We showcase the effectiveness of our MoTCoder in addressing challenging code generation tasks, specifically focusing on the following prominent benchmarks:
APPS (Hendrycks et al., 2021b) is a description-to-code generation benchmark from competitive programming platforms Codewars [^3] https://www.codewars.com/, AtCoder [^4] https://atcoder.jp/, Kattis [^5] https://open.kattis.com/, Codeforces [^6] https://codeforces.com/, etc. Building upon prior research (Hendrycks et al., 2021b; Chen et al., 2021c; Li et al., 2022a), we conducted an assessment of the models using the passing rate metric pass@k. This metric is defined as the proportion of problems successfully solved by employing k generated programs for each problem.
CodeContests (Li et al., 2022a) is a competitive programming dataset sourced from Aizu [^7] https://judge.u-aizu.ac.jp, AtCoder [^8] https://atcoder.jp, CodeChef [^9] https://www.codechef.com, Codeforces, HackerEarth [^10] https://www.hackerearth.com, etc. Building upon prior research (Hendrycks et al., 2021b; Chen et al., 2021c; Li et al., 2022a), we conducted an assessment of the models using the passing rate metric pass@k.
4.1 Examination of Modular Ability
In this section, we evaluate the modular capabilities of our model through both normal and modular inference, utilizing prompts as illustrated in Figure 5. For modular prompts, we instruct the model to utilize provided submodules to complete the final code. Conventional models lack this ability to leverage submodules, making the logical and modular aspects more challenging for the model.
Results are in Table 1. On one hand, our MoTCoder demonstrates a significant performance boost compared to models with the same parameters trained with traditional approaches, achieving a remarkable 12.9% improvement on APPs. On the other hand, when employing modular reasoning on models that have not undergone MoT training, there is a considerable decrease in performance. In contrast, our outcomes are minimally affected. This confirms the superior modular capabilities of our approach, which is crucial in large-scale code development environments, where developers often leverage and integrate code sub-modules to construct code frameworks, rather than crafting each line of code from scratch. MoTCoder enables the construction of complete code using only a standardized description of functions, unlocking the potential to generate extensive and complex code frameworks.
The results of MoTCoder also demonstrates that the absence of submodule inputs has a negligible impact on the results. This implies that our method autonomously learns to design submodules without requiring additional hints. In subsequent experiments, we use normal prompts during the inference process for MoTCoder and other methods to ensure a fair comparison.
4.2 Ablation Experiments
In this section, we perform ablation experiments to examine the effectiveness of our MoTCoder training framework.
4.2.1 Modular-of-Thought Instruction-Tuning
We propose two instruction-tuning strategies:
  1. 1.
    One-step strategy. The models directly generate the necessary modules and incorporate them into a complete solution.
  2. 2.
    Two-step strategy. Initially, the instructions outline the essential sub-modules, generating their function headers and docstrings that describe their intended purpose. Subsequently, the instruction guides the model to implement these outlined modules and eventually merge them into a cohesive final solution.
Note that both strategies generate substeps within a single process. As indicated in Table 2, the suggested two-step modular-of-thought instruction tuning strategy outperforms the other both for normal and modular inference. This aligns with the problem-solving approach employed by adept developers: When faced with a problem, skilled developers design solutions with a modular structure, organizing them into high-level logical sub-modules. They subsequently finalize and integrate these components to effectively produce their final solutions.
4.2.2 Training Dataset
In this section, we conduct ablation experiments on the dataset, as illustrated in Table 3. Initially, we finetuned our model on the modular dataset obtained after modular-of-though instruction evolution, resulting in a substantial improvement (11.0%) compared to the baseline of normal fine-tuning. Subsequently, we explore the incorporation of a normal dataset, conducting experiments on a mixture of normal and modular datasets, revealing further performance enhancement of 3.38%. This underscores the significance of data diversity: While we encourage the model to decompose tasks into submodules, we also allow the model to employ direct methods to solve problems. This provides a larger solution space, leading to significant improvements, especially for easy and medium-level problems (2.46% and 4.54%, respectively). The improvement is more modest for difficult problems (0.84%), aligning with the intuition that direct methods are better suited for simpler problems.
Table 6: CodeContests test and validation results by pass@k (%). The rounds of revision after the first round of direct generation is 5 for CodeChain (Le et al., 2023). SOTA models includes code-davinci, GPT3.5 and WizardCoder (Luo et al., 2023b), with code-revision method CodeChain (Le et al., 2023) and filtering method CoT(Chen et al., 2023).
Model Size Revision Validation Test Average
pass@1 pass@5 pass@1 pass@5 pass@1 pass@5
code-davinci-002 - - - - 1.00 - 1.00 -
code-davinci-002 + CodeT - - - - 3.20 - 3.20 -
WizardCoder 15B - 1.11 3.18 1.98 3.27 1.55 3.23
WizardCoder + CodeChain 15B 5 2.35 3.29 2.48 3.30 2.42 3.30
MoTCoder 15B - 2.39 7.69 6.18 12.73 4.29 10.21
GPT3.5 - - 6.81 16.23 5.82 11.16 6.32 13.70
4.3 Results
In this section, we present a comprehensive comparison of our approach with existing large coding model baselines on APPS  (Hendrycks et al., 2021a) in Section 4.3.1 and CodeContest (Li et al., 2022a) in Section 4.3.2. We employ normal inference prompts for all methods.
4.3.1 Results on APPS
We conducted a comparison of our approach with existing large language model baselines on APPs (Hendrycks et al., 2021a). All outcomes are computed using raw predictions without being filtered by the test cases provided in the prompt. Our analysis includes a comparison with open-sourced approaches such as CodeT5 (Wang et al., 2021), fine-tuned GPT-Neo (Hendrycks et al., 2021a), GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), one-shot StarCoder (Li et al., 2023), and one-shot WizardCoder (Luo et al., 2023b). Additionally, we present results from closed-source models, including text-davinci, code-davinci, and GPT3.5.
As depicted in the results from the APPS shown in Table 4, we notice that the modular-of-thought training approach leads to an enhancement in code generation capabilities compared with previous instruction finetuning models. Our MoTCoder exhibits improved performance across all difficulty levels and demonstrates more substantial gains in the interview and competition-level problems, characterized by more intricate solutions. To provide specifics, the pass@1 performance of MoTCoder-15B surpasses WizardCoder-15b by 15.49% on interview-level problems and 10.28% on competition-level problems. Moreover, our pass@1 on competition-level questions surprisingly outperformed the closed-source model GPT3.5 by an impressive margin of 5.7%. This outcome serves as compelling evidence of the efficacy of our approach in addressing intricate problems.
We also conduct a comparative analysis of our approach against previous LLM baselines with code-revision methods as well. Our included baselines contain Codex (Chen et al., 2021c), CodeT5 (Wang et al., 2021), code-davinci, StarCoder (Li et al., 2023), and WizardCoder (Luo et al., 2023b) and code-revision methods contain Self-edit (Zhang et al., 2023), CodeRL (Wang et al., 2021; Le et al., 2022), Self-repair (Olausson et al., 2023), and CodeChain (Le et al., 2023). The results presented in Table 5 illustrate that MoTCoder exhibits notable performance improvements.
Specifically, MoTCoder achieves a 20.80% pass@1, surpassing the current SOTA, CodeChain + WizardCoder, by 10.3%. CodeChain (Le et al., 2023) is an iterative framework for inference that elicits modularized code generation through a chain of self-revisions. In contrast to CodeChain, our approach introduces a guided modular-of-thought framework during training, making it more intrinsic. This leads to improvements in performance. Notably, this achievement is obtained without the need for any additional revision iterations, thereby avoiding any extra associated inference costs.
4.3.2 Results on CodeContests
We conduct a comprehensive evaluation of our approach on CodeContests, benchmarking it against current state-of-the-art models, including code-davinci, GPT3.5, and WizardCoder (Luo et al., 2023b). The competitive models utilize the code-revision method CodeChain (Le et al., 2023) and the filtering method CodeT (Chen et al., 2023). The results, as depicted in Table 6, reveal notable performance enhancements achieved by MoTCoder. Specifically, MoTCoder attains an impressive 12.73% pass@5 on the test set, surpassing WizardCoder with CodeChain by 9.43%. Remarkably, MoTCoder even outperforms GPT3.5 by 1.57%. Importantly, these achievements are realized without the need for additional revision iterations or filtering by test samples, thereby avoiding any additional associated inference costs or the introduction of extra information.
5 Limitations
Through experimentation, it has been observed that although our MoTCoder elevates the model’s modular and logical capabilities, leading to a commendable performance in addressing competitive-level programming challenges, it concurrently sacrifices some of the model’s conversational abilities. This finding, along with corroborative evidence from prior research (Roziere et al., 2023), underscores the persistent and crucial nature of addressing the forgetting problem. We remain committed to ongoing exploration of this issue in our future work.
6 Conclusion
We propose a novel framework, MoTCoder, for tuning instructions based on a modular of thought. MoTCoder directs LLMs to decompose their solutions into modular segments, culminating in the synthesis of a comprehensive final solution. Our experiments demonstrate that modular-of-thought strategy significantly improves LLM performance, achieving SOTA passing rate on challenging code tasks and showcasing superior modular capabilities. It is crucial in large-scale code development environments, since developers commonly leverage code sub-modules to construct code frameworks over crafting each line of code from scratch. MoTCoder empowers the creation of entire code structures using only standardized descriptions of functions. Adopting this methodology, we unlock the potential to generate extensive and intricate code frameworks.