### v3

# MoTCoder: Elevating Large Language Models with Module-of-Thought

Abstract
Large Language Models (LLMs) have showcased impressive capabilities in handling straightforward programming tasks. However, their performance tends to falter when confronted with more challenging programming problems. We observe that conventional models often generate solutions as monolithic code blocks, restricting their effectiveness in tackling intricate questions. To overcome this limitation, we present Module-of-Thought Coder (MoTCoder). We introduce a framework for MoT instruction tuning, designed to promote the decomposition of tasks into logical sub-tasks and sub-modules. Our investigations reveal that, through the cultivation and utilization of sub-modules, MoTCoder significantly improves both the modularity and correctness of the generated solutions, leading to substantial relative pass@1 improvements of 60% on APPS and 19% on CodeContests. Our codes are available at https://github.com/dvlab-research/MoTCoder.
1 Introduction
Developing systems that can generate executable and functionally correct computer programs has long been sought after in artificial intelligence (Manna & Waldinger, 1971). Recently, Large Language Models (LLMs) (Brown et al., 2020; OpenAI, 2023a; Chowdhery et al., 2022; Anil et al., 2023; Hoffmann et al., 2022; Rae et al., 2021; Zeng et al., 2022; Touvron et al., 2023; Zhang et al., 2022) have showcased remarkable success in many problem domains beyond natural language processing, and is poised as a promising approach to tackle code modeling and generation (as well as other coding related tasks) (Rozière et al., 2023; Black et al., 2021; Chen et al., 2021c). Through instruction fine-tuning (Li et al., 2023b; Luo et al., 2023b; Wang et al., 2023a; Li et al., 2022b; Nijkamp et al., 2023; Zheng et al., 2023; Fried et al., 2022; Chen et al., 2021a; Wang et al., 2021), LLMs have achieved impressive performance on code generation benchmarks like HumanEval (Chen et al., 2021b) and MBPP (Austin et al., 2021).
Yet, when confronted with intricate coding problems such as APPS (Hendrycks et al., 2021b) and CodeContests  (Li et al., 2022a), SOTA models struggle to match seasoned developers (Hendrycks et al., 2021b; Li et al., 2022a; Shinn et al., 2023). The main culprit is their overly simplistic generation approach. Current models produce the code solution as a single monolithic block in an attempt to solve the problem in one shot. While feasible for simple tasks, this approach becomes increasingly inefficient to tackle a complex task which naturally entails solving multiple sub-tasks. In contrast, adept developers often devise modularized solutions by breaking down the original problem into more approachable components that can be individually and more efficiently solved.
Following this intuition, recent works (Jiang et al., 2023; Le et al., 2023) propose iterative code inference for code generation. They, however, come with added inference costs and offer only marginal performance gains. In this work, we propose to more efficiently improve the modularization capability of coding LLMs using Module-of-Thought (MoT) instruction fine-tuning. Our approach guides LLMs to break down their solution into modular components, each representing an abstract function dedicated to a logical sub-task. To train the model to adhere to the MoT prompt, we generate instructional data using a process termed MoT Code Instruction Transformation. In this process, LLMs are instructed to outline necessary modules, generating only their function headers and docstrings that describe their intended usage. Subsequently, the instruction guides the model to implement these modules and eventually combine them into the final solution. After that, we fine-tune the LLM with our MoT instruction fune-tuning, resulting in our MoTCoder model.
In summary, our contributions are threefold:
  1. 1.
    We propose a 2-step Module-of-Thought Code Instruction Transformation approach that instructs LLMs to generate modularized code solutions, where each module is to address a logical sub-task and accompanied by standardized annotations.
  2. 2.
    We develop Module-of-Thought Coder (MoTCoder), a new model that enhances the modularization capabilities of LLMs with MoT Instruction Tuning. It breaks down complex problems into manageable, logically coherent sub-modules, improving both comprehensibility and maintainability of the generated code.
  3. 3.
    Our experiments demonstrate that MoTCoder establishes new SOTA results on challenging code generation benchmarks such as APPS and CodeContests. Specifically, MoTCoder improves the pass@1 performance by significant margins, exemplified by improvements over existing metrics by 60% on APPS and 19% on CodeContests, as illustrated in Section 1.
[…]
figurePass@1 results on CodeContests (left) and APPS (right). Comparison of our MoTCoder with previous SOTAs.
2 Related Works
2.1 Large Language Models
General LLMs. In recent times, Large Language Models (LLMs) have exhibited remarkable prowess across a wide array of tasks. Leading technology companies have made significant advancements in developing highly proficient closed-source LLMs, including OpenAI’s GPT3 (Brown et al., 2020) and GPT4 (OpenAI, 2023a), Google’s PaLM (Chowdhery et al., 2022; Anil et al., 2023), Bard [^1] https://bard.google.com/, DeepMind’s Chinchilla (Hoffmann et al., 2022), and Gopher (Rae et al., 2021), as well as Anthropic’s Claude [^2] https://www.anthropic.com/index/introducing-claude. The AI community has also observed the release of several open-source LLMs, where model weights are made publicly available. EleutherAI has contributed GPT-NeoX-20B (Black et al., 2022) and GPT-J-6B (Wang & Komatsuzaki, 2021). Google has released UL2-20B (Tay et al., 2022). Tsinghua University has introduced GLM-130B (Zeng et al., 2022). Meta has released OPT (Zhang et al., 2022) and LLaMA (Touvron et al., 2023).
Coding LLMs. Recent research has introduced a significant number of LLMs tailored for code-related tasks to address the challenges of code understanding and generation. Closed-source models include OpenAI’s Codex (Chen et al., 2021a) and Code-Davinci (Microsoft, 2023). Google has proposed PaLM-Coder (Chowdhery et al., 2022). These models excel on popular code completion benchmarks such as HumanEval (Chen et al., 2021b) and MBPP (Austin et al., 2021). On the open-source front, Salesforce has introduced CodeGen (Nijkamp et al., 2023), CodeT5 (Wang et al., 2021), and CodeT5+ (Wang et al., 2023b). Tsinghua University has contributed CodeGeeX (Zheng et al., 2023), and the BigCode Project has developed StarCoder (Li et al., 2023b). These models have demonstrated significant advancements in code-related tasks.
2.2 Instruction Fine-Tuning
General Instruction Tuning. In its early stages, the core aim of instruction fine-tuning was to amplify the cross-task generalization capabilities of Language Models (LMs). This was accomplished by subjecting LMs to fine-tuning using an extensive corpus of public Natural Language Processing (NLP) tasks. Pioneering this approach, T5 (Raffel et al., 2020) underwent training on a diverse set of supervised text-to-text tasks. Subsequent endeavors like FLAN (Wei et al., 2022a), ExT5 (Aribandi et al., 2022), T0 (Sanh et al., 2022a), and UnifiedQA (Khashabi et al., 2020) broadened the spectrum of tasks, fortifying the overall generalization capability of LMs. Noteworthy contributions from ZeroPrompt (Xu et al., 2022) and FLAN-T5 (Chung et al., 2022) pushed boundaries by incorporating thousands of tasks into their training pipelines. OpenAI has taken an alternative route by enlisting human annotators to contribute an extensive corpus of human instructions, encompassing diverse formats and a broad spectrum of task types. Building upon this dataset, OpenAI trained its GPT-3 (Brown et al., 2020) model to create InstructGPT (Ouyang et al., 2022), which better aligns with users’ inputs. This developmental trajectory has given rise to notable works such as ChatGPT. In the open-source realm, Alpaca (Taori et al., 2023) adopts the self-instruct method (Wang et al., 2022), leveraging ChatGPT to generate data for training. Vicuna (Chiang et al., 2023) utilizes user-shared conversations collected from ShareGPT.com to train its models. Introducing the Evol-Instruct method, WizardLM (Xu et al., 2023) involves evolving existing instruction data to generate more intricate and diverse datasets.
Chain-of-Thought Instruction Tuning. In contrast to these generalized approaches, Recent works (Luo et al., 2023c; Yue et al., 2023; Chen et al., 2022b; Gunasekar et al., 2023; Haluptzok et al., 2023) apply instruction tuning method in common-sense reasoning (West et al., 2022), text-summarization (Sclar et al., 2022), mathematical reasoning  (Luo et al., 2023a; Yue et al., 2023), tool use (Patil et al., 2023), coding (Luo et al., 2023b), and general-purpose reasoning (Li et al., 2023a; Zelikman et al., 2022). Among them,  Yue et al. (2023) curates a diverse corpus of mathematics problems with chain-of-thought or program-of-thought (Chen et al., 2022b) annotations for mathematical reasoning analogous to our plans. Gunasekar et al. (2023) proposed pre-training models on programming textbooks generated synthetically from GPT3.5. Haluptzok et al. (2023) similarly generates programming puzzles and corresponding solutions from language models.
2.3 Prompting Techniques
A prompt technique known as Chain of Thought (CoT) (Wei et al., 2022b) generated intermediate reasoning steps before the answer to address language reasoning tasks. Inspired by CoT, a series of prompting works has been proposed. Least-to-most prompting (Zhou et al., 2022) reduces a complex problem into a series of sub-problems and then solves the sub-problems in order, adding the answer to the previous sub-problems to the prompt each time solving begins. PAL (Gao et al., 2022) and PoT (Chen et al., 2022a) is proposed to generating code as the intermediate reasoning steps, delegating solving to the compiler, thus improving solution accuracy. Similar methods are proposed for simple mathematical (Lewkowycz et al., 2022; Wu et al., 2022), commonsense (Sanh et al., 2022b; Madaan et al., 2022), symbolic reasoning (Yao et al., 2023) and code generation problems (Jiang et al., 2023; Le et al., 2023). However, these works can only plan code during generation. In comparison, our approach introduces a guided module-of-thought framework during training, making it more intrinsic. Our investigations reveal that, through the cultivation and utilization of sub-modules, MoTCoder significantly enhances both the modularity and correctness of the generated solutions.
Figure 1: Left: To transform conventional instructions into module-of-thought instructions, our MoT instruction transformation approach guides the models through a two-step process. Initially, the models are instructed to outline the necessary sub-modules, generating only their function headers and docstrings that describe their intended usage. The subsequent instruction then guides the model to implement these sub-modules and integrate them into a cohesive final solution. This instruction is complemented by a one-shot example to further enhance understanding. Right: Examples of instruction, modularized code and solution.
3 Methods
In this section, we first detail the module-of-thought instruction transformation in Section 3.1 and then introduce the module-of-thought instruction tuning strategy for our MoTCoder in Section 3.2.
3.1 Module-of-Thought Instruction Transformation
In this section, we first introduce the normal instruction, followed by our module-of-thought instruction and its assessment.
Figure 2: Illustration of our 2-step Module-of-Thought Instruction Transformation: Plain code utilizes a single module to directly generate code. In contrast, our Module-of-Thought Instruction Transformation first instructs LLMs to outline necessary sub-modules, generating only their function headers and docstrings that describe their intended usage. Subsequently, the instructions guide the model in implementing these sub-modules and eventually combining them into a comprehensive final solution.
3.1.1 Normal Instruction
In general, a code sequence generated by a language model $\theta$ through the autoregressive sampling of tokens $\hat{o}_{t}$ is from the parameterized conditional distribution:
$o_{t}\sim p_{\theta}(.|o_{1:t-1},I),$ (1)
where $I$ represents the input instruction and $o_{t}$ is the $t$-th token of the flattened output sequence.
3.1.2 Module-of-Thought Instruction
Considering programming encompasses both detailed coding abilities and broader logical or strategic planning capabilities, we propose a transformation pipeline that enhances code clarity and structure targeting detailed coding proficiencies and integrating descriptions based on natural language for the broader strategic planning capabilities.
According to previous researches (Jain et al., 2023), straightforward and unambiguous, detailed instructions enhance the model’s efficacy and precision in executing the desired tasks. Hence, for intricate data transformation tasks, decomposing the task into simpler, sequential steps yields better outcomes. Therefore, our proposed methodology aims to transform normal instructions into a sequential code generation process by leading the models through a two-step procedure, as illustrated in Figure 2 .
  1. 1.
    Sub-modules. Initially, the models are instructed to outline the required sub-modules, generating only their function headers and docstrings describing their intended usage.
    $\hat{S_{i}}\sim p_{\theta}(.|\hat{S}_{1:i-1},I),$ (2)
    where $\hat{S}_{i}$ represents the $i$-th sub-module outlined by the model and $I$ represents the input instruction.
  2. 2.
    Final solution. The subsequent instruction guides the model to implement these sub-modules and eventually combine them into a comprehensive final solution.
    $\hat{o}_{t}\sim p_{\theta}(.|\hat{o}_{1:t-1},\{\hat{S}_{i}\},I),$ (3)
    where $\hat{o}_{t}$ is the $t$-th of the flattened output sequence.
The instruction is supplemented with a one-shot example, serving to prompt the model to adhere to the MoT instruction generation strategy. An illustration of the instruction prompt is presented in Figure 1. The instruction encourages the model to decompose a program into sub-modules. This mirrors the methodology commonly employed by developers when addressing intricate coding tasks, where they systematically break down solutions into modular components.
3.1.3 Instruction Assessment
Throughout these transformations, we introduce guidelines for reviewing our transformed module-of-thought code. We identify the situations below as markers of instruction refinement failure:
  1. 1.
    The refined instruction diverges from the module-of-thought generation strategy, not adhering to the protocol of initial sub-module creation followed by the main code development.
  2. 2.
    At the sub-module creation phase, if no sub-modules are formed or if overarching code is developed instead.
  3. 3.
    During the main code development phase, the absence of main code creation or the emergence of multiple main code blocks indicates a problem.
  4. 4.
    The presence of test cases in the dataset that the transformed program fails to pass. This criterion ensures the transformed programs preserve functional equivalence with the original codes.
3.2 Module-of-Thought Instruction Tuning
In this section, we first introduce our module-of-thought dataset, followed by our MoT instruction tuning details.
3.2.1 MoT Dataset
We use data sourced from a small part of the training set of APPS (Hendrycks et al., 2021b) and CodeContests (Li et al., 2022a) as our initial set. For each problem, we take at most one answer, whereas the original dataset may contain dozens of answers. We then use GPT3.5 to generate transformed instructions, with the instruction transformation prompt as shown in Figure 1. The generated module-of-thought instruction dataset comprises around 24k instructions. We merge the docstrings that describe the intended usage and input-output variables type of the functions with their implementation. Some examples from the MoT dataset are in Appendix B.
3.2.2 MoT Instruction Tuning
Our MoTCoder underwent instruction tuning utilizing the SOTA coding model Deepseek-Cdoer-Base (Balog et al., 2016) across three epochs on our proposed MoT instruction dataset. The model’s maximum input length is 2048 tokens. We used a training batch size of 16 and an evaluation batch size of 1 per device, with gradient accumulation steps set to 4. The learning rate was initialized at $2\times 10^{-5}$, with a warmup period of 10 steps to gradually adapt the learning rate, following a cosine learning rate scheduler for optimization. Additionally, we enabled gradient checkpointing, BF16 precision and more deepseed settings to enhance training efficiency and reduce memory consumption: Our model was configured with the AdamW optimizer and utilized the WarmupLR scheduler to manage the learning rate adjustments effectively. To optimize memory and compute resources, we employed a third-stage zero optimization setting and enable communication overlap, contiguous gradients, and large sub group size of $1\times 10^{9}$ to streamline the training process. In addition, the maximum live parameters, maximum reuse distance, and the parameter settings for gathering 16-bit weights during model saving were all set to $1\times 10^{9}$. This process took approximately 28 hours using 8 A800 GPUs.
4 Experiments
We demonstrate the efficacy of MoTCoder in tackling intricate code-generation tasks, rather than those that can be solved with just a few lines, as exemplified by HumanEval (Chen et al., 2021b) and MBPP (Austin et al., 2021). Specifically, we focus on the following prominent benchmarks.
APPS (Hendrycks et al., 2021b) is a description-to-code generation benchmark from competitive programming platforms Codewars [^3] https://www.codewars.com/, AtCoder [^4] https://atcoder.jp/, Kattis [^5] https://open.kattis.com/, Codeforces [^6] https://codeforces.com/, etc. Building upon prior research (Hendrycks et al., 2021b; Chen et al., 2021c; Li et al., 2022a), we conducted an assessment of the models using the passing rate metric pass@k. This metric is defined as the proportion of problems successfully solved by employing k generated programs for each problem.
CodeContests (Li et al., 2022a) is a competitive programming dataset sourced from Aizu [^7] https://judge.u-aizu.ac.jp, AtCoder [^8] https://atcoder.jp, CodeChef [^9] https://www.codechef.com, Codeforces, HackerEarth [^10] https://www.hackerearth.com, etc. Building upon prior research (Hendrycks et al., 2021b; Chen et al., 2021c; Li et al., 2022a), we conducted an assessment of the models using the passing rate metric pass@k.
Table 1: Training Datasets. We compare MoTCoder training on normal and modular-of-thought dataset. APPs test results by pass@1 (%) are shown.
Model Introductory Interview Competition All
Normal 50.20 26.65 11.52 28.33
Module-of-Thought 50.01 29.81 14.36 30.76
Table 2: APPS test results by pass@k (%). Competitive approaches include CodeT5 (Wang et al., 2021), fine-tuned GPT-Neo (Hendrycks et al., 2021a), GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), one-shot StarCoder (Li et al., 2023b), WizardCoder (Luo et al., 2023c), CodeLlama series (Roziere et al., 2023), text-davinci, GPT-4 (OpenAI, 2023b). We also include code-revision methods containing Self-edit (Zhang et al., 2023), CodeRL (Le et al., 2022), Self-repair (Olausson et al., 2023), and CodeChain (Le et al., 2023).
Model Size Introductory Interview Competition All
CodeT5 770M 6.60 1.03 0.30 2.00
CodeRL+CodeT5 770M 7.08 1.86 0.75 2.69
text-davinci-002 - - - - 7.48
Self-edit+text-davinci-002 - - - - 7.94
GPT-2 0.1B 5.64 6.93 4.37 6.16
1.5B 7.40 9.11 5.05 7.96
GPT-Neo 2.7B 14.68 9.85 6.54 10.15
GPT-3 175B 0.57 0.65 0.21 0.55
StarCoder 15B 7.25 6.89 4.08 6.40
WizardCoder 15B 26.04 4.21 0.81 7.90
CodeChain+WizardCoder 15B 26.29 7.49 3.75 10.50
Octocoder 16B 16.50 7.92 4.61 8.97
Codellama 7B 14.15 6.63 4.00 7.61
13B 23.94 13.50 9.80 14.85
34B 32.01 18.61 10.19 19.61
Codellama-Python 7B 18.83 8.62 4.47 9.83
13B 26.40 13.44 6.86 14.72
34B 26.45 16.61 8.77 17.01
Codellama-Instruct 7B 14.20 6.63 4.43 7.70
13B 22.41 14.34 6.62 15.21
34B 28.64 16.80 10.51 17.91
Deepseek-Coder-Base 6.7B 40.23 22.12 13.04 23.92
Deepseek-Coder-Instruct 6.7B 44.65 23.86 12.89 25.83
MoTCoder 6.7B 50.01 29.81 14.36 30.76
GPT-4 - 34.97 13.75 14.52 18.15
4.1 Ablation Experiments
In this section, we conduct ablation experiments on the dataset, as illustrated in Table 1. We evaluated the performance of instruction tuning on our proposed Module-of-Thought (MoT) dataset versus original dataset, which is a small part of training set of CodeContests and APPS. We utilize the DeepSeek-Coder-6.7B model as the base model. The results, as delineated in the table below, highlight a pivotal finding: while the MoT tuning strategy maintains comparable performance on introductory tasks, it significantly enhances outcomes in competitive programming scenarios, elevating the pass@1 rate from 11.52% to 14.36%. The results align with our intuition that breaking down code into sub-modules is more suitable for complex tasks, whereas simpler tasks are better suited for direct generation. The overall enhancement underlines the effectiveness of our MoT method in elevating model proficiency in competitive programming domains.
Table 3: CodeContests test results by pass@k (%). The rounds of revision after the first round of direct generation is 5 for CodeChain Le et al. (2023). SOTA models include StarCoder Li et al. (2023b), WizardCoder Luo et al. (2023c), CodeLlama series models Roziere et al. (2023), code-davinci, GPT4 OpenAI (2023b), with code-revision method CodeChain Le et al. (2023) and filtering method CoTChen et al. (2023).
Model Size pass@1 pass@5
code-davinci-002 - 1.00 -
code-davinci-002 + CodeT - 3.20 -
WizardCoder 15B 1.98 3.27
WizardCoder + CodeChain 15B 2.48 3.30
Octocoder 16B 4.95 13.03
Codellama 7B 0.30 1.11
13B 2.12 6.26
34B 5.35 12.02
Codellama-Python 7B 4.75 10.30
13B 4.75 12.32
34B 5.86 14.85
Codellama-Instruct 7B 2.12 6.26
13B 5.96 12.02
34B 6.46 14.24
Deepseek-Coder-Base 6.7B 6.46 15.25
Deepseek-Coder-Instruct 6.7B 6.87 8.18
MoTCoder 6.7B 9.29 16.97
GPT-4 - 16.36 -
4.2 Results
In this section, we present a comprehensive comparison with existing large coding model baselines on APPS  (Hendrycks et al., 2021a) and CodeContest (Li et al., 2022a).
4.2.1 Results on APPS
We conducted a comparison of our approach with existing large language model baselines on APPs (Hendrycks et al., 2021a). All outcomes are computed using raw predictions without being filtered by the test cases provided in the prompt. Our analysis includes a comparison with open-sourced approaches such as CodeT5 (Wang et al., 2021), fine-tuned GPT-Neo (Hendrycks et al., 2021a), GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), one-shot StarCoder Li et al. (2023b), WizardCoder (Luo et al., 2023c), and CodeLlama series (Roziere et al., 2023). Additionally, we present results from SOTA closed-source models GPT-4 (OpenAI, 2023b).
As depicted in the results from the APPS shown in Table 2, we notice that the module-of-thought training approach leads to an enhancement in code generation capabilities compared with previous instruction finetuning models. Our MoTCoder exhibits improved performance across all difficulty levels and demonstrates more substantial gains in the interview and competition-level problems, characterized by more intricate solutions. To provide specifics, the pass@1 performance of MoTCoder-6.7B surprisingly outperformed the closed-source model GPT-4 by an impressive margin of 12.61%. This outcome serves as compelling evidence of the efficacy of our approach in addressing competitive programming problems.
We also conduct a comparative analysis of our approach against previous LLM baselines with code-revision methods as well. Our included baselines contain Codex (Chen et al., 2021c), CodeT5 (Wang et al., 2021), code-davinci, StarCoder (Li et al., 2023b), and WizardCoder (Luo et al., 2023c) and code-revision methods contain Self-edit (Zhang et al., 2023), CodeRL (Wang et al., 2021; Le et al., 2022), Self-repair (Olausson et al., 2023), and CodeChain (Le et al., 2023). The results presented in Table 2 illustrate that MoTCoder exhibits notable performance improvements.
Specifically, MoTCoder achieves a 30.76% pass@1, surpassing CodeChain + WizardCoder, by 20.26%. CodeChain (Le et al., 2023) is an iterative framework for inference that elicits modularized code generation through a chain of self-revisions. In contrast to CodeChain, our approach introduces a guided module-of-thought framework during training, making it more intrinsic. This leads to improvements in performance. Notably, this achievement is obtained without the need for any additional revision iterations, thereby avoiding any extra inference costs.
4.2.2 Results on CodeContests
We conduct an evaluation of our approach on CodeContests (Li et al., 2022a), benchmarking it against current state-of-the-art models, including code-davinci, GPT3.5, and WizardCoder (Luo et al., 2023c). The competitive models utilize the code-revision method CodeChain (Le et al., 2023) and the filtering method CodeT (Chen et al., 2023). The results, as depicted in Table 3, reveal notable performance enhancements achieved by MoTCoder. Specifically, MoTCoder attains an impressive 12.73% pass@5 on the test set, surpassing WizardCoder with CodeChain by 9.43%. These achievements are realized without the need for additional revision iterations or filtering by test samples, thereby avoiding any additional associated inference costs or the introduction of extra information.
5 Limitations
Through experimentation, it has been observed that although our MoTCoder elevates the model’s modular and logical capabilities, leading to a commendable performance in addressing competitive-level programming challenges, it concurrently sacrifices some of the model’s conversational abilities. This finding, along with corroborative evidence from prior research Roziere et al. (2023), underscores the persistent and crucial nature of addressing the forgetting problem. We remain committed to ongoing exploration of this issue in our future work.
6 Conclusion
This study highlights the limitations of Large Language Models (LLMs) in solving complex programming tasks due to their tendency to generate monolithic code blocks. In response, we developed Module-of-Thought Coder (MoTCoder), a novel framework that encourages the breakdown of tasks into manageable sub-tasks and sub-modules. Our results demonstrate that MoTCoder’s approach significantly enhances the modularity and accuracy of solutions, as evidenced by considerable improvements in pass@1 rates on both APPS and CodeContests benchmarks. We believe the introduction of MoT instruction tuning as a method to cultivate and leverage sub-modules paves the path for a promising direction for future research.
Appendix A Broader impacts
Positive societal impacts.
We propose a method for challenging code programming, which will be helpful for users to solve code problems.
negative societal impacts.
This approach may be abused in programming contests that are supposed to be done by humans.
Appendix B Examples
For given input problems, we provided submodules and final code solution examples for problems from APPs test dataset.
B.0.1 APPs Example 1
Listing 1: Problem 1 from APPs test dataset
An accordion is a string (yes, in the real world accordions are musical instruments, but lets forget about it for a while) which can be represented as a concatenation of: an opening bracket (ASCII code \$091\$), a colon (ASCII code \$058\$), some (possibly zero) vertical line characters (ASCII code \$124\$), another colon, and a closing bracket (ASCII code \$093\$). The length of the accordion is the number of characters in it.
For example, [::], [:||:] and [:|||:] are accordions having length \$4\$, \$6\$ and \$7\$. (:|:), {:||:}, [:], ]:||:[ are not accordions.
You are given a string \$s\$. You want to transform it into an accordion by removing some (possibly zero) characters from it. Note that you may not insert new characters or reorder existing ones. Is it possible to obtain an accordion by removing characters from \$s\$, and if so, what is the maximum possible length of the result?
—–Input—–
The only line contains one string \$s\$ (\$1 $\le$ |s| $\le$ 500000\$). It consists of lowercase Latin letters and characters [, ], : and |.
—–Output—–
If it is not possible to obtain an accordion by removing some characters from \$s\$, print \$-1\$. Otherwise print maximum possible length of the resulting accordion.
—–Examples—–
Input
|[a:b:|]
Output
4
Input
Output
-1
Listing 2: Submodules generated for Problem 1 from APPs test dataset
def find_colon_indices(s):
Description: This function finds the indices of the first and last occurrence of ":" in the given string.
Input:
s (str): A string containing colons.
Output:
ind (int): The index of the first occurrence of ":".
bind (int): The index of the last occurrence of ":".
def count_pipes_between_colons(s, ind, bind):
Description: This function counts the number of "|" characters between the first and last occurrence of ":" in the given string.
Input:
s (str): A string containing colons and "|" characters.
ind (int): The index of the first occurrence of ":".
bind (int): The index of the last occurrence of ":".
Output:
ans (int): The count of "|" characters between the first and last occurrence of ":".
Listing 3: Code solution generated for Problem 1 from APPs test dataset
def find_colon_indices(s):
ind = s.find(":")
bind = s.rfind(":")
return ind, bind
def count_pipes_between_colons(s, ind, bind):
ans = 0
if ind!= 0:
ans += s.count("|", 0, ind)
if bind!= len(s) - 1:
ans += s.count("|", bind + 1, len(s))
return ans
s = input()
ind, bind = find_colon_indices(s)
ans = count_pipes_between_colons(s, ind, bind)
if ind == -1 or bind == -1 or ind > bind:
print(-1)
else:
print(bind - ind + 1 + ans)
B.0.2 APPs Example 2
Listing 4: Problem 2 from APPs test dataset
Anton has the integer x. He is interested what positive integer, which doesnt exceed x, has the maximum sum of digits.
Your task is to help Anton and to find the integer that interests him. If there are several such integers, determine the biggest of them.
—–Input—–
The first line contains the positive integer x \$(1 $\leq$ x $\leq$ 10^18)\$ - the integer which Anton has.
—–Output—–
Print the positive integer which doesnt exceed x and has the maximum sum of digits. If there are several such integers, print the biggest of them. Printed integer must not contain leading zeros.
—–Examples—–
Input
100
Output
99
Input
48
Output
48
Input
521
Output
499
Listing 5: Submodules generated for Problem 2 from APPs test dataset
def find_best_number(num):
Description: This function finds the best number by decrementing digits from the given number.
Input:
num (list): A list of digits representing the number.
Output:
best (list): A list of digits representing the best number.
def convert_to_string(num):
Description: This function converts a list of digits into a string representation.
Input:
num (list): A list of digits.
Output:
s (str): The string representation of the digits.
Listing 6: Code solution generated for Problem 2 from APPs test dataset
def find_best_number(num):
best = num[:]
for i in range(8, -1, -1):
if num[i] > 0:
num[i] -= 1
break
return num
def convert_to_string(num):
s = .join(map(str, num))
return s
x = int(input())
num = [int(i) for i in str(x)]
best = find_best_number(num)
s = convert_to_string(best)
print(s)
print(s)